OpenStudy (anonymous):

PLEASE HELP!!! 1/2 x + 1/3 y = 7 and 1/4 x + 2/3 y = 6 Y = 5, X = 10 2/3 Why?

OpenStudy (anonymous):

hi. thanks for coming over to help =)

OpenStudy (johnweldon1993):

so first....turn them into slope intercept form...that is....isolate y

OpenStudy (anonymous):

ok...working that on paper right now...

OpenStudy (anonymous):

(1/2x + 1/3y = 7)3...1 1/2x + y = 21...y = 21 - 1 1/2 x...correct?

OpenStudy (johnweldon1993):

(1/2x + 1/3y = 7) FIRST isolate the y number....so subtract the 1/2 x from both sides...what do you have now?

OpenStudy (anonymous):

you have 1/3y = 7 - 1/2x

OpenStudy (johnweldon1993):

correct...now...to isolate y ...how do we get rid of the 1/3? we multiply by 3 correct? so do that to both sides...now what do you have?

OpenStudy (anonymous):

y = 21 - 3/2x ... so y = 21 - 1 1/2x

OpenStudy (johnweldon1993):

right...but keep it as an improper fraction...it's easier than a mixed number but now for the next one 1/4 x + 2/3 y = 6 using the same process...what does this come out to?

OpenStudy (anonymous):

solving for x this time?

OpenStudy (johnweldon1993):

No you're again isolating y...you'll see why in the next step after you find it

OpenStudy (anonymous):

ok...doing right now...

OpenStudy (anonymous):

1/4x + 2/3 y = 6...2/3y = 6 - 1/4x...y = (6 - 1/4x) 2/3...

OpenStudy (anonymous):

* i mean times 3/2 at the end

OpenStudy (anonymous):

then y = 18/2 -3/8x

OpenStudy (johnweldon1993):

correct.....just simplify totally...so y = 9 -3/8x

OpenStudy (johnweldon1993):

so now what 2 equations do you have?

OpenStudy (anonymous):

i have y = 9 -3/8 and y = 21 - 3/2x

OpenStudy (anonymous):

*-3/8x

OpenStudy (anonymous):

so now i would combine them, correct?

OpenStudy (johnweldon1993):

right..the x after the -3/8x so...now put them in proper form...y=mx+b......slope in front y intercept at the end

OpenStudy (johnweldon1993):

just rewrite them for clarity.... y = -3/8x + 9 y = -3/2x + 21 And yes...now combine them

OpenStudy (johnweldon1993):

that is...set them equal to each other

OpenStudy (anonymous):

2y = 18/16 + 30

OpenStudy (anonymous):

then i divide both sides by 2...

OpenStudy (johnweldon1993):

no no... i shouldnt have said combine them like i did.....that's why i corrected your statement....what do you do when y = 2 different things?

OpenStudy (anonymous):

that's y = 9/8 + 15

OpenStudy (johnweldon1993):

say you have y = 2 and y= 3 that is the same thing as saying y = 2 AND 3 right? in that same logic....that must mean 2 = 3 so in THIS tense....is y = both of those...set them equal to each other

OpenStudy (anonymous):

that means that something is strange with the equation? like, if i were to graph it, it would be different shape than a line?

OpenStudy (johnweldon1993):

meaning do it like this -3/8x + 9 = -3/2x + 21

OpenStudy (anonymous):

ok. now what?

OpenStudy (johnweldon1993):

now you can see that the only variable you have is x right? that's what we're solving for.....so combine your x's and combine your constant numbers...then isolate x......what do you have?

OpenStudy (johnweldon1993):

but do you understand that logic behind setting the 2 equations equal to each other?

OpenStudy (anonymous):

not exactly...but I did the equation: 26/16x = 18...right?

OpenStudy (johnweldon1993):

what is the least common denominator of -3/8x and -3/2x?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

wait no- 8!

OpenStudy (johnweldon1993):

....correct with the 8......what do you need to multiply 2 by to get 8?

OpenStudy (anonymous):

4

OpenStudy (johnweldon1993):

correct....so multiply 3/2 by 4/4

OpenStudy (anonymous):

3/4

OpenStudy (johnweldon1993):

alright...lets lead up to what you've done so far -3/8x + 9 = -3/2x + 21 subtract 9 from both sides -3/8x = -3/2x + 12 now add -3/2x to both sides.... -3/8x + 3/2x = 12 this is where your least common denominator comes in..which is 8 (-3/8x) + (4/4) times 3/2x = 12 so -3/8x + 12/8x = 12

OpenStudy (johnweldon1993):

combine those x's...what do we have?

OpenStudy (anonymous):

alright i was just rereading what we did...9/8x = 12

OpenStudy (johnweldon1993):

perfect.....now how do you isolate x here?

OpenStudy (anonymous):

multiply both sides by 8/9

OpenStudy (johnweldon1993):

correct...and what does that come to?

OpenStudy (anonymous):

x = 96/9, or x = 10 and 6/9, which can be reduced to x = 10 and 2/3

OpenStudy (anonymous):

yay we did X =)

OpenStudy (johnweldon1993):

exactly! that's where your x comes from.....now for y....how would you find it? now that you have x?

OpenStudy (anonymous):

I would substitute what x equalls and take an equation to find y... 1/2 (10 and 2/3) + 1/3 y = 7

OpenStudy (johnweldon1993):

that is correct

OpenStudy (johnweldon1993):

and so what do we get for y?

OpenStudy (anonymous):

that would be 5 and 1/3 + 1/3y = 7...then 1/3y = 1 and 2/3...then...y = 3 and 1/2

OpenStudy (anonymous):

that's not right....

OpenStudy (johnweldon1993):

Ahh i see what you did..... so 1/2 x + 1/3 y = 7 1/2 (what we got for x) + 1/3 y = 7 ***turn the mixed number into an improper fraction 1/2(32/3) + 1/3 y = 7 so yes this is 5 and 1/3 7 - 5 and 1/3 IS 1 and 2/3 turn 1 and 2/3 into an improper fraction..what is that?

OpenStudy (anonymous):

5/3

OpenStudy (johnweldon1993):

correct...so we have 1/3y = 5/3 how do we isolate y?

OpenStudy (anonymous):

multiply by 3 y = 15/3...so... y = 5 YAY!!!

OpenStudy (johnweldon1993):

haha...and that is your reasoning :)

OpenStudy (anonymous):

thank you so much! i don't know what i'd do without open study and helpful people like you. have a great day! or evening...or whatever time it is on your side of the world... =)

OpenStudy (johnweldon1993):

lol you're welcome...glad to help!